With a total flow of 0.4 m3/s, determine the division of flow and the head loss from A to B.

With a total flow of 0.4 m3/s, determine the divis
Answer #1

1)Total flow Q = flow in upper pipe Q1+ flow in in lower pipe Q2

2) Flow Q = Area of pipe A * velocity of flow V

From 1 and 2 ,

Q = A1 * v1 + A2 *v2

Q = \pi * D_{1}^{2} / 4 *v_{1} + \pi * D_{2}^{2} / 4 *v_{2}

Q = (\pi /4) *( D_{1}^{2} *v_{1} + D_{2}^{2}*v_{2})

0.4 = (\pi /4) *( 0.55^{2} *v_{1} + 0.3^{2}*v_{2})

0.5093= ( 0.55^{2} *v_{1} + 0.3^{2}*v_{2})

0.5093= ( 0.3025 *v_{1} +0.09 * v_{2}) ------------ Equation A

3) The head loss in each branch pipe WILL BE THE SAME!

hL = h1 = h2

4) h_{L} = f * (L/D) * v^{2}/2g

From 3 and 4

f_{1} * (L_{1}/D_{1}) * v_{1}^{2}/2g =f_{2} * (L_{2}/D_{2}) * v_{2}^{2}/2g

0.018* (1800/0.55) * v_{1}^{2}/2g =0.018 * (600/0.3) * v_{2}^{2}/2g

Cancelling 2g on both sides

58.91 * v_{1}^{2}=36 * v_{2}^{2}

v_{1}}=0.7817 * v_{2} ----- Equation B

Solving equation A and B

0.5093=0.3025 *0.7817 * v_{2} + 0.09* v_{2}

v_{2} =1.56 m/s

Then, v_{1} =0.7817 *1.56

v_{1} =1.22 m/s

Flow in upper pipe Q1 = A1 * v1 = \pi * 0.55^2 /4 * 1.22 = 0.289 m3/ s

Q2 = A2 * v2 =\pi * 0.3^2 /4 * 1.56 = 0.11 m3/s

Head loss from A to B =

0.018 * (1800/0.55) *1.22^{2}/(2 *9.8) = 4.47 m